Write a program equivalent to the C language assignment statement
z = a + (b * c) + (d * e) - (f / g) - (h * i);
Use only:
(a) Three-address instructions
(b) Two-address instructions

Please explain the process as clearly as possible. I would like to LEARN how this is done in so little address instructions. Thank you.
Answer #1

z = a + (b * c) + (d * e) - (f / g) - (h * i)

(a) Three-address instructions

It is assumed that the computer has two processor registers, R1 and R2. The symbol M [A] denotes the operand at memory address A.

Every instruction has 2 source operands and a destination register/memory address

Instruction format : Operation destination operand, source operand1, source operand2.

destination operand = source operand1 Operation source operand2.

MUL R1, b, c -------- R1 <- M[b] * M[c] R1 = b * c

MUL R2, d, e -------- R2 <- M[d] * M[e] R2 = d * e

ADD R1, R1, R2 ---- R1 <- R1 + R2 R1 = (b * c) + (d * e)

DIV R2, f, g ---------- R2 <- M[f] / M[g] R2 = f / g

SUB R1, R1, R2 ---- R1 <- R1 - R2 R1 = (b * c) + (d * e) - (f / g)

MUL R2, h, i --------- R2 <- M[h] * M[i] R2 = h * i

SUB R1, R1, R2 ---- R1 <- R1 - R2 R1 = (b * c) + (d * e) - (f / g) - (h * i)

ADD z, R1, R2 ------ z <- R1 + R2   z = a + (b * c) + (d * e) - (f / g) - (h * i)

(b) Two-address instructions

Every instruction has 2 source operands and the first source operand acts as the destination register/memory address

Instruction format : Operation destination operand, source operand

destination operand = destination operand Operation source operand.

MOV R1, b ------ R1 <- M[b] R1 = b

MUL R1, c ------ R1 <- R1 * c R1 = b*c

MOV R2, d ----- R2 <- M[d] R2 = d

MUL R2, e ------ R2 <- R2 * e R2 = d*e

ADD R1, R2 ---- R1 <- R1 + R2 R1 = (b * c) + (d * e)

MOV R2, f ------ R2 <- M[f] R2 = f

DIV R2, g ------- R2 <- R2 / g R2 = f/g

SUB R1, R2 ---- R1 <- R1 - R2 R1 = (b * c) + (d * e) - (f / g)

MOV R2, h ----- R2 <- M[h] R2 = f

MUL R2, i ------ R2 <- R2 * i R2 = h*i

SUB R1, R1, R2 ----- R1 <- R1 - R2 R1 = (b * c) + (d * e) - (f / g) - (h * i)

ADD R1, a -------- R1 <- R1 + a R1 = a + (b * c) + (d * e) - (f / g) - (h * i)

MOV z, R1 ------- z <- R1 z = a + (b * c) + (d * e) - (f / g) - (h * i)